OpenStudy (anonymous):

Solve for x: 3(x+2)^1/2 = 24

OpenStudy (anonymous):

i got one question the ^meant multiply to a power ?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

i got u hold on

OpenStudy (anonymous):

x = 7

OpenStudy (anonymous):

\[\sqrt{x+2}=24/3 x+2=24^{2}/3^{2} x+2=64 x=62

OpenStudy (anonymous):

@joyful312 @nocturn4l

OpenStudy (anonymous):

x=62

OpenStudy (anonymous):

are u asking or correcting

OpenStudy (anonymous):

thanks to both of you either way but i think @Hanishala has the answer

OpenStudy (anonymous):

correcting

OpenStudy (anonymous):

wat how is it 62 when nun of those number were in the equation. like im stuck

OpenStudy (anonymous):

use 7 in the equation you will find that 3*sqrt(7 )=24 and its false but 3* sqrt(62+2)=3 *sqrt(64) =3*8 =24 and its correct

OpenStudy (anonymous):

it didnt say nun about sqrt

OpenStudy (anonymous):

but i guess

OpenStudy (anonymous):

(x+2)^1/2 = sqrt(x+2) @nocturn4l

OpenStudy (anonymous):

the question was 3(x+2)^1/2 no sqrt. im still stuck but i tried

OpenStudy (anonymous):

(x+2)^1/2 = sqrt(x+2)

OpenStudy (anonymous):

im done

OpenStudy (anonymous):

LMAO! @joyful312 im crying

OpenStudy (anonymous):

"im done" me in math class

OpenStudy (anonymous):

heyyy im serious lol. but thatts exactly how i was in AP U.S History

OpenStudy (anonymous):

Had to give you a mdal for that

OpenStudy (anonymous):

lol ok ok . i guess i couldnt take the fact that i was wrong

OpenStudy (anonymous):

is it wrong that x=62 @Alexandra675

OpenStudy (anonymous):

no one knows we all crappy

OpenStudy (anonymous):

i wouldve said the other word but they might kick me off for cursing

OpenStudy (anonymous):

i am sure that x=62

OpenStudy (anonymous):

it doesn't @Hanishala stop using wolfram alpha

OpenStudy (anonymous):

wait so i was right

OpenStudy (anonymous):

this is the equation he used \[3\sqrt{x+2}=24\]

OpenStudy (anonymous):

now tell me what that equals

OpenStudy (anonymous):

62 right?

OpenStudy (anonymous):

yeah but was i right

OpenStudy (anonymous):

idk tho

OpenStudy (anonymous):

man im right im claiming it

OpenStudy (anonymous):

lol u got the medal anyways

OpenStudy (anonymous):

that dont mean im right i probably got it because u think im funny

OpenStudy (anonymous):

lol